substitution : Y=2X-4
X+Y=-3

Answer by rapaljer(4551)

Y=2X-4
X+Y=-3

Take the "y=___" equation and substitute this in place of the y in the second equation:
x+(_____)=-3
x+(2x-4) =-3
3x-4=-3
3x-4+4=-3+4
3x=1
x=1/3

Now, substitute this value of x back into the "y=" equation.
y=2x-4
y=2(1/3)-4
y=2/3 -4

Get a common denominator of 3, so change -4 to -12/4:
y=2/3-12/3
y=-10/3

Check in the second equation:
x+y=-3
1/3+(-10/3)=-3
-9/3=-3 It checks!

For additional explanation on this topic, please go to my website. Do a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" at the top of this search. Look near the top of my Homepage for "Basic, Intermediate and College Algebra: One Step at a Time", click on "Basic Algebra" and look in Chapter 4 for Section 4.07 Solving Systems of Equations.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus